# Hodge structures on conformal blocks

### v1

@@ 3.4. The Cohomological Field Theory of a Hodge structure
The main reference for the definition of a CohFT is [20, 0.5].
@@ 3.5. Polarizations of Hodge structures on modular functors
A similar definition can be provided for complex rooted functors.
@@ 4.1. The $\mathrm{SU}(2)$ modular functors
For each $r\geq 3$ odd integer, the $\mathrm{SU}(2)$ modular functor of level $2r$ is defined over the cyclotomic field $\mathbb{Q}(\zeta_{4r})$, has set of colors $\Lambda=\{0,1,2,\dotsc,r-2\}$ and has a hermitian structure $h$ (coming from the Skein construction). All of its quantum representations are semisimple, as for $\zeta_{r}:=\zeta_{4r}^{4}=e^{(r-1)i\pi/r}$, $h$ is unitary. Moreover, whenever $\sum_{i}\lambda_{i}$ is odd, for all $g$, $\mathcal{V}_{g}(\underline{\lambda})=0$.
The following theorem gives a geometric construction of the rooted functor associated to $\mathcal{V}_{\zeta_{r}}$. The notation $\mathcal{V}^{\mathrm{ab}}_{\zeta_{r}}$ is for the abelian modular functor of level $2r$, see [16, 2.1.3].
Theorem 4.2 ([16, 3.3]).
@@
From this fact and computations we proved that gluing maps are compatible with these geometric Hodge structures ([16, 3.18]). This provides the following.
@@
Let $r\geq 3$ be odd. The rooted functor associated to $\mathcal{V}_{\zeta_{r}}$ supports a rational Hodge structure such that each $\mathcal{V}_{\zeta_{r}}(b;a_{1},\dotsc,a_{n})$ has weight $m=\frac{a_{1}+\dotsb+a_{n}-b}{2}$ and all Hodge bidegrees $(p,q)$ are pairs of non-negative integers.
@@
In their recent paper [7], Belkale, Fakhruddin and Mukhopadhyay independently computed the same rings $V$ by computing coefficients $a_{k}$ and $b_{k}$ (defined below) with a different method (see [7, Th. 1.11]). See Section 4.2 below for more comments on their work.
@@ 4.2. Modular functors associated to simple Lie Algebras
Proposition 4.8.
For any $\mathfrak{g}$ and $\zeta$ as above, the modular functor $\mathcal{V}_{\mathfrak{g},\zeta}$ has semisimple quantum representations and its associated weakly ribbon category is ribbon.
Proof.
The functors $\mathcal{V}_{\mathfrak{g},\zeta}$ for different choices of $\zeta$ are Galois-conjugate and for $\zeta=e^{\frac{i\pi}{2m(l+h^{\vee})}}$, its quantum representations are unitary (see [6], the genus 00 case implies the general case). This explains semisimplicity. For the fact that the associated category is ribbon, see [4, 3.3]. ∎
@@
In [7], for any $\mathfrak{g}$ and $\zeta$ as above, Belkale, Fakhruddin and Mukhopadhyay study the Hodge theory of a geometrically constructed framed 999see [16, 2.3] for the definition of the framing of a rooted functor. rooted functor $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ whose quantum representations coincide with that of $\mathcal{V}_{\mathfrak{g},\zeta}$. In particular, they provide an algorithm for computing the Hodge numbers when $\mathfrak{g}$ is $\mathfrak{sl}_{n}$.
@@
The framed rooted functor associated to $\mathcal{V}_{\mathfrak{g},\zeta}$ is isomorphic to $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$. Equivalently, $\mathcal{V}_{\mathfrak{g},\zeta}$ and $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ have isomorphic braided tensor categories.
To prove this Conjecture, essentially, one needs to identify gluing isomorphisms of both framed rooted functors. Assuming the Conjecture, we can, as in the case of $\mathrm{SU}(2)$ modular functors, apply unicity of the Hodge structure to get the following result.
@@
Let $n\geq 2$, $\ell\geq 1$ and $\zeta$ be a $2(\ell+n)$-th root of unity. Then if $\mathcal{V}_{\mathfrak{sl}_{n},\zeta}$ satisfies 4.10, the algorithm [7, 1.11] of Belkale, Fakhruddin and Mukhopadhyay computes all Hodge numbers of $\mathcal{V}_{\mathfrak{sl}_{n},\zeta}$ (for any genus $g$).
@@ 4.3. Gaps in Hodge decompositions
For simplicity this section is written for modular functors. But, with the exception of Proposition 4.18, all definitions and statements extend to genus 00 modular functors and rooted functors.
@@
It is known that at prime and $2$ times prime levels, the corresponding $\mathrm{SO}(3)$/$\mathrm{SU}(2)$ quantum representations are all irreducible (for $n=0$ this is [21], for a proof in general, see for example the discussion in [15, A.]). However, at more general levels, little is known about irreducibility. Corollary 4.16 tells us that Hodge numbers will be of no help in investigating this question.
@@
The assumptions of Proposition 4.18 are satisfied by any $\mathrm{SU}(2)$ or $\mathrm{SO}(3)$ modular functor.
5. The weakly ribbon category of a modular functor
5.1. Description of the weakly ribbon category
@@
We will abbreviate $\mathrm{Vect}_{\Lambda}^{R}$ to $\mathrm{Vect}_{\Lambda}$ when $R$ is obvious. For definitions of monoidal categories, rigid monoidal categories, braided monoidal categories, and twists on braided monoidal categories, see [12, 2.1, 2.10, 8.1 and 8.10]. Note that in this section we assume all structures to be linear but do not assume rigidity.
Let us now describe how to construct from a rooted functor $\mathcal{V}$ a braided monoidal structure with a twist on $\mathrm{Vect}_{\Lambda}$. Because of how $\mathrm{Vect}_{\Lambda}$ is defined, we only need to define the structure on the generating set of simple objects $\{[\lambda]\mid\lambda\in\Lambda\}$. We define the monoidal unit by $1=[0]$, the monoidal product by:
@@
The braiding $\beta_{\lambda,\mu}:[\lambda]\otimes[\mu]\rightarrow[\mu]\otimes[\lambda]$ is given on the $[\nu]$ components by the action of the braid $\beta_{1}$ on $\mathcal{V}(D^{2},\nu;\lambda,\mu)$. As for the twist $\theta_{\lambda}:[\lambda]\rightarrow[\lambda]$, it is given, after the canonical identification $[\lambda]=R[\lambda]\simeq\mathcal{V}(D^{1},\lambda;\lambda)[\lambda]$, by the action of the Dehn twist $\theta_{1}$ on $\mathcal{V}(D^{1},\lambda;\lambda)$. We can also simply describe $\theta_{\lambda}$ as multiplication by $t_{\lambda}$. As this is quite folklore, we leave it to the reader to check that the axioms of a braided monoidal structure with a twist are satisfied (otherwise, an equivalent but slightly different definition of the structures, together with a proof, are provided in [4, 5.3]).
Let us remark that because of axiom (Dual) in Definition 2.14, the monoidal structure on $\mathrm{Vect}_{\Lambda}$ is weakly rigid in the sense of [4, 5.3.4], i.e. for each object $x$, the functor $y\mapsto\hom(1,x\otimes y)$ is (left) representable by an object $x^{*}$, called its weak right dual, and the functor $x\mapsto x^{*}$ is an equivalence. In our case, the weak duals are given on simple objects by $[\lambda]^{*}=\mathcal{V}(D^{2},0;\lambda,\lambda^{\dagger})^{\vee}[\lambda^{\dagger}]$. We shall use Bakalov and Kirillov’s convention and call a weakly rigid braided monoidal category with a twist a weakly ribbon category.
@@
Let $\mathcal{V}$ be a rooted functor over $R$ with set of colors $\Lambda$. Then $\mathcal{V}$ induces on $\mathrm{Vect}_{\Lambda}$ the structure of a weakly ribbon category. We shall denote this category with structure $\mathrm{C}_{\mathcal{V}}$.
@@
We know define a groupoid of weakly ribbon categories. Note that our notion of isomorphism is non-standard.
@@
For $R$ a ring, let $\mathfrak{wRib}$ be the category whose:
objects are pairs $(\Lambda,\mathrm{Vect}_{\Lambda})$, where $\Lambda$ is some finite set, and $\mathrm{Vect}_{\Lambda}$ is endowed with a weakly ribbon structure;
morphisms from $(\Lambda_{1},\mathrm{Vect}_{\Lambda_{1}})$ to $(\Lambda_{2},\mathrm{Vect}_{\Lambda_{2}})$ are pairs $(f,\phi)$, with $f:\Lambda_{1}\rightarrow\Lambda_{2}$ a bijection and $\phi:\otimes_{1}\simeq f^{*}\otimes_{2}$ a natural isomorphism, such that they induce a monoidal isomorphism $\mathrm{Vect}_{\Lambda_{1}}\rightarrow\mathrm{Vect}_{\Lambda_{2}}$ compatible with the braidings and twists. In other words, morphisms form $(\Lambda_{1},\mathrm{Vect}_{\Lambda_{1}})$ to $(\Lambda_{2},\mathrm{Vect}_{\Lambda_{2}})$ are ribbon isomorphisms that are induced by a bijection $f:\Lambda_{1}\rightarrow\Lambda_{2}$ at the level of the $R$-linear categories and are strict on tensor units  [^10] Strictness of $f^{*}[0]=[0]\stackrel{{\scriptstyle\mathrm{id}}}{{\simeq}}[0]$ is a consequence of the normalization axiom of rooted functors..
@@ 5.2. Full-faithfulness of forgetful functors
Let $R$ be a ring and $c\in R^{\times}$. Denote by $\mathfrak{tmod}_{c}$ the full subcategory of $T\mathfrak{mod}$ consisting of modular functors with central charge $c$. Then all forgetful functors in the sequence:
$\mathfrak{tmod}_{c}\longrightarrow\mathfrak{mod}^{0}\longrightarrow\mathfrak{root}\longrightarrow\mathfrak{wRib}$
@@
By Proposition 5.4, the map $\mathfrak{root}\rightarrow\mathfrak{wRib}$ is faithful. Moreover, as any surface can be cut into genus 00 surfaces, we see that $\mathfrak{mod}_{c}\rightarrow\mathfrak{mod}^{0}$ is also faithful. Let us turn to the functor $\mathfrak{mod}^{0}\rightarrow\mathfrak{root}$. For any colored genus 00 surface $\Sigma=(S,\underline{\lambda})$ with at least one boundary component $B\subset\partial S$, say numbered $1$, we have for any $\mathcal{V}\in\mathfrak{mod}^{0}$, by definition:
@@
canonically. As the right-hand side depends only on the rooted functor associated to $\mathcal{V}$, $\mathfrak{mod}^{0}\rightarrow\mathfrak{root}$ is faithful. ∎
The aim of the rest of this section is to prove fullness. Note that by faithfulness, we need only prove it for $\mathfrak{tmod}_{c}\rightarrow\mathfrak{wRib}$.
@@
We expect Theorem 5.6 to be a consequence of a much more general result of Brochier and Woike [9, 6.8]. However, their definition of the category of modular functors is more complicated and non-explicit. Some work would be necessary to simplify their definition in the case of semisimple modular functors and compare it to ours. In this paper, we provide a proof that does not use their results.

### v2

@@ 3.4. The Cohomological Field Theory of a Hodge structure
The main reference for the definition of a CohFT is [27, 0.5].
@@ 3.5. Polarizations of Hodge structures on modular functors
A similar definition can be provided for complex ribbon and braided functors.
@@ 4.1. The $\mathrm{SU}(2)$ modular functors
For each $r\geq 3$ odd integer, the $\mathrm{SU}(2)$ modular functor of level $2r$ is defined over the cyclotomic field $\mathbb{Q}(\zeta_{4r})$, has set of colors $\Lambda=\{0,1,2,\dotsc,r-2\}$ and has a Hermitian structure $h$ (coming from the Skein construction). Moreover, whenever $\sum_{i}\lambda_{i}$ is odd, for all $g$, $\mathcal{V}_{g}(\underline{\lambda})=0$.
The following theorem gives a geometric construction of the braided functor associated to $\mathcal{V}_{\zeta_{r}}$. The notation $\mathcal{V}^{\mathrm{ab}}_{\zeta_{r}}$ is for the Abelian modular functor of level $2r$, see [18, 2.1.3].
Theorem 4.2 ([18, 3.3]).
@@
From this fact and computations we proved that gluing maps are compatible with these geometric Hodge structures ([18, 3.18]). This provides the following.
@@
Let $r\geq 3$ be odd. The braided functor associated to $\mathcal{V}_{\zeta_{r}}$ supports a rational Hodge structure such that each $\mathcal{V}_{\zeta_{r}}(b;a_{1},\dotsc,a_{n})$ has weight $m=\frac{a_{1}+\dotsb+a_{n}-b}{2}$ and all Hodge bidegrees $(p,q)$ are pairs of non-negative integers.
@@
In their recent paper [6], Belkale, Fakhruddin and Mukhopadhyay independently computed the same rings $V$ by computing coefficients $a_{k}$ and $b_{k}$ (defined below) with a different method (see [6, Th. 1.11]). See Section 4.2 below for more comments on their work.
@@ 4.2. Modular functors associated to simple Lie Algebras
In [6], for any $\mathfrak{g}$ and $\zeta$ as above, Belkale, Fakhruddin and Mukhopadhyay study the Hodge theory of a geometrically constructed braided functor $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ whose quantum representations coincide with those of $\mathcal{V}_{\mathfrak{g},\zeta}$. In particular, they provide an algorithm for computing the Hodge numbers when $\mathfrak{g}$ is $\mathfrak{sl}_{n}$.
@@
The braided functor associated to $\mathcal{V}_{\mathfrak{g},\zeta}$ is isomorphic to $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$. Equivalently, $\mathcal{V}_{\mathfrak{g},\zeta}$ and $\widetilde{\mathcal{V}}_{\mathfrak{g},\zeta}$ have isomorphic braided tensor categories.
To prove this Conjecture, essentially, one needs to identify gluing isomorphisms of both braided functors. Assuming the Conjecture, we can, as in the case of $\mathrm{SU}(2)$ modular functors, apply unicity of the Hodge structure to get the following result.
@@
Let $n\geq 2$, $\ell\geq 1$ and $\zeta$ be a $2(\ell+n)$-th root of unity. Then if $\mathcal{V}_{\mathfrak{sl}_{n},\zeta}$ satisfies 4.9, the algorithm [6, 1.11] of Belkale, Fakhruddin and Mukhopadhyay computes all Hodge numbers of $\mathcal{V}_{\mathfrak{sl}_{n},\zeta}$ (for any genus $g$).
@@ 4.3. Gaps in Hodge decompositions
For simplicity this section is written for modular functors. But, with the exception of Proposition 4.17, all definitions and statements extend to genus 00 modular functors, ribbon and braided functors.
@@
It is known that at prime and $2$ times prime levels, the corresponding $\mathrm{SO}(3)$/$\mathrm{SU}(2)$ quantum representations are all irreducible (for $n=0$ this is [28], for a proof in general, see for example the discussion in [17, A.]). However, at more general levels, little is known about irreducibility. Corollary 4.15 tells us that Hodge numbers will be of no help in investigating this question.
@@
The assumptions of Proposition 4.17 are satisfied by any $\mathrm{SO}(3)$ modular functor.
5. The ribbon fusion category of a modular functor
5.1. Description of the ribbon fusion category
@@
We will abbreviate $\mathrm{Vect}_{\Lambda}^{R}$ to $\mathrm{Vect}_{\Lambda}$ when $R$ is obvious. For definitions of monoidal category, rigid monoidal category, fusion category, braided monoidal category, ribbon tensor category and modular category, see [12, 2.1, 2.10, 4.1, 8.1, 8.10 and 8.13].
Let us now describe how to construct from a ribbon functor $\mathcal{V}$ a ribbon structure on $\mathrm{Vect}_{\Lambda}$, and from a braided functor $\mathcal{V}$ braided structure on $\mathrm{Vect}_{\Lambda}$. Because of how $\mathrm{Vect}_{\Lambda}$ is defined, we only need to define the structure on the generating set of simple objects $\{[\lambda]\mid\lambda\in\Lambda\}$. We define the monoidal unit by $1=[0]$, the monoidal product by:
@@
The braiding $\beta_{\lambda,\mu}:[\lambda]\otimes[\mu]\rightarrow[\mu]\otimes[\lambda]$ is given on the $[\nu]$ components by the action of the braid $\beta_{1}$ on $\mathcal{V}(D^{2},\nu;\lambda,\mu)$. As for the twist $\theta_{\lambda}:[\lambda]\rightarrow[\lambda]$ in the case of ribbon functors, it is given, after the canonical identification $[\lambda]=R[\lambda]\simeq\mathcal{V}(D^{1},\lambda;\lambda)[\lambda]$, by the action of the Dehn twist $\theta_{1}$ on $\mathcal{V}(D^{1},\lambda;\lambda)$. We can also simply describe $\theta_{\lambda}$ as multiplication by $t_{\lambda}$. As this is quite folklore, we leave it to the reader to check that the axioms of a braided monoidal structure (with a twist) are satisfied (otherwise, an equivalent but slightly different definition of the structures, together with a proof, are provided in [4, 5.3]).
Let us remark that because of axiom (Dual) in Definition 2.14, the monoidal structure on $\mathrm{Vect}_{\Lambda}$ is weakly rigid in the sense of [4, 5.3.4], i.e. for each object $x$, the functor $y\mapsto\hom(1,x\otimes y)$ is (left) representable by an object $x^{*}$, called its weak right dual, and the functor $x\mapsto x^{*}$ is an equivalence. In our case, the weak duals are given on simple objects by $[\lambda]^{*}=\mathcal{V}(D^{2},0;\lambda,\lambda^{\dagger})^{\vee}[\lambda^{\dagger}]$.
Moreover, Etingof and Penneys recently proved the long-standing conjecture that braided weakly fusion categories are braided fusion categories [14]. Hence the category $(\mathrm{Vect}_{\Lambda},\otimes,c,\beta)$ is braided fusion, and in the case where $\mathcal{V}$ is a ribbon functor, $(\mathrm{Vect}_{\Lambda},\otimes,c,\beta,\theta)$ is ribbon.
@@
Let $\mathcal{V}$ be a braided, respectively ribbon, functor over $R$ with set of colors $\Lambda$. Then $\mathcal{V}$ induces on $\mathrm{Vect}_{\Lambda}$ the structure of a braided fusion category, respectively ribbon fusion category. We shall denote this category with structure $\mathrm{C}_{\mathcal{V}}$.
@@
We know define a groupoids of ribbon fusion categories and braided fusion categories. Note that our notion of isomorphism is non-standard.
@@
For $R$ a ring, let $\mathfrak{Rib}$ be the category whose:
objects are pairs $(\Lambda,\mathrm{Vect}_{\Lambda})$, where $\Lambda$ is some finite set, and $\mathrm{Vect}_{\Lambda}$ is endowed with a ribbon structure;
morphisms from $(\Lambda_{1},\mathrm{Vect}_{\Lambda_{1}})$ to $(\Lambda_{2},\mathrm{Vect}_{\Lambda_{2}})$ are pairs $(f,\phi)$, with $f:\Lambda_{1}\rightarrow\Lambda_{2}$ a bijection and $\phi:\otimes_{1}\simeq f^{*}\otimes_{2}$ a natural isomorphism, such that they induce a monoidal isomorphism $\mathrm{Vect}_{\Lambda_{1}}\rightarrow\mathrm{Vect}_{\Lambda_{2}}$ compatible with the braidings and twists. In other words, morphisms form $(\Lambda_{1},\mathrm{Vect}_{\Lambda_{1}})$ to $(\Lambda_{2},\mathrm{Vect}_{\Lambda_{2}})$ are ribbon isomorphisms that are induced by a bijection $f:\Lambda_{1}\rightarrow\Lambda_{2}$ at the level of the $R$-linear categories and are strict on tensor units  [^9] Strictness of $f^{*}[0]=[0]\stackrel{{\scriptstyle\mathrm{id}}}{{\simeq}}[0]$ is a consequence of the normalization axiom of ribbon functors..
One similarly defines the category $\mathfrak{Braid}$ of braided fusion categories.
@@ 5.2. Full-faithfulness of forgetful functors
Let $R$ be a ring, all of whose quotients have characteristic 00 and $c\in R^{\times}$. Denote by $\mathfrak{mod}_{c}$ the full subcategory of $\mathfrak{mod}$ consisting of modular functors with central charge $c$. Then all horizontal [^10] The vertical ones are certainly not full. forgetful functors in the commutative diagram:
[…]
@@
By Proposition 5.4, the functors $\mathfrak{ribfun}\rightarrow\mathfrak{Rib}$ and $\mathfrak{braidfun}\rightarrow\mathfrak{Braid}$ are faithful. Moreover, as any surface can be cut into genus 00 surfaces, we see that $\mathfrak{mod}_{c}\rightarrow\mathfrak{mod}^{0}$ is also faithful. Let us turn to the functor $\mathfrak{mod}^{0}\rightarrow\mathfrak{ribfun}$. For any colored genus 00 surface $\Sigma=(S,\underline{\lambda})$ with at least one boundary component $B\subset\partial S$, say numbered $1$, we have for any $\mathcal{V}\in\mathfrak{mod}^{0}$, by definition:
@@
canonically. As the right-hand side depends only on the ribbon functor associated to $\mathcal{V}$, $\mathfrak{mod}^{0}\rightarrow\mathfrak{ribfun}$ is faithful. ∎
The aim of the rest of this section is to prove the fullnesses.
@@
We expect Theorem 5.6 to be a consequence of a much more general result of Brochier and Woike [8, 6.8]. However, their definition of the category of modular functors is more complicated and non-explicit. Some work would be necessary to simplify their definition in the case of semisimple modular functors and compare it to ours. In this paper, we provide a proof that does not use their results.
5.3. Remarks on equivalences of functors with categories and on truncations
In this subsection, we explain why, over $\mathbb{C}$, the fully-faithful functors in Theorem 5.6 are (almost) equivalences, and also extend the Theorem to the case of truncated functors. We will need neither of these facts in this paper, but place them here to give the reader a fuller picture and for future reference.
Notation 5.8.
We will denote by $\mathfrak{Mod}\subset\mathfrak{Rib}$ the full sub-groupoid of modular fusion category, i.e. those whose $S$-matrix is non-degenerate (see [12, 8.13.4]).
Theorem 5.9.
Here, the base ring is $R=\mathbb{C}$. Consider the commutative diagram of forgetful functors
[…]
Then $\mathrm{f}^{0}$, $\mathrm{f}^{r}$ and $\mathrm{f}^{b}$ are equivalences of groupoids, each $\mathrm{f}_{c}$ is fully-faithful, and all objects of $\mathfrak{Mod}$ are in the essential image of $\mathrm{f}_{c}$ for some $c\in\mathbb{C}^{\times}$  [^11] See [4, 5.7.10] for the exact values $c$ may take..
Proof.
The facts about the essential images of $\mathrm{f}^{r}\circ\mathrm{f}^{0}$ and $\bigsqcup_{c}\mathrm{f}_{c}$ are proved in [4, 5.4.1, 6.7.13]. For $\bigsqcup_{c}\mathrm{f}_{c}$, this is the Reshetikhin-Turaev construction. All full-faithfulnesses are part of Theorem 5.6. These imply that $\mathrm{f}^{0}$ and $\mathrm{f}^{r}$ are equivalences. Note that with our definition a braided functor is equivalent to an $E_{2}$-algebra in the monoidal $2$-category of $\mathbb{C}$-linear semisimple categories. Then, the fact that $\mathrm{f}^{b}$ is essential surjective follows from the known fact that $E_{2}$-algebras in the category of categories are equivalent to braided monoidal categories, see [24, 1.2.4, 1.2.5]. ∎
Let us turn to the case of truncated functors. These are obtained by restricting the possible Euler characteristic of surfaces in the definitions.
Definition 5.10.
For $\chi\leq-3$ an integer, denote by $\mathrm{S\tilde{u}rf}_{\Lambda}^{\geq\chi}\subset\mathrm{S\tilde{u}rf}_{\Lambda}$ the full subgroupoid of surfaces of Euler characteristic $\geq\chi$. By replacing $\mathrm{S\tilde{u}rf}_{\Lambda}$ with $\mathrm{S\tilde{u}rf}_{\Lambda}^{\geq\chi}$ in the Definition 2.37 of modular functor, one obtains the definition of $\chi$-truncated modular functor. The groupoid of $\chi$-truncated modular functors with central charge $c$ will be denoted $\mathfrak{mod}_{c}^{\geq\chi}$.
One similarly defines $\chi$-truncated genus 00 modular functors, ribbon functors and braided functors. The corresponding groupoids will be denoted $\mathfrak{mod}^{0,\geq\chi}$, $\mathfrak{ribfun}^{\geq\chi}$ and $\mathfrak{braidfun}^{\geq\chi}$ respectively.
Theorem 5.11.
Here, the base ring is $R=\mathbb{C}$. Fix $c\in\mathbb{C}^{\times}$. Then for any $\chi\leq-3$, the following forgetful functors are well-defined and equivalences:
$\mathfrak{mod}_{c}\longrightarrow\mathfrak{mod}_{c}^{\geq\chi}\longrightarrow\mathfrak{Mod}.$
The same result holds for genus 00 modular functors, ribbon functors and braided functors, replacing $\mathfrak{Mod}$ by $\mathfrak{Rib}$, $\mathfrak{Rib}$ and $\mathfrak{Braid}$ respectively.
Remark 5.12.
Note that $\chi\leq-3$ is required for $\mathfrak{mod}_{c}^{\geq\chi}\longrightarrow\mathfrak{Rib}$ to be defined.
Proof.
The non-trivial part of the well-definiteness is that categories in the essential image of $\mathfrak{mod}_{c}^{\geq\chi}\rightarrow\mathfrak{Rib}$ are modular. When Bakalov-Kirillov prove in [4, 5.5.1] that the ribbon fusion category associated to a modular functor is modular, they rely on relations in tori with $1$ and $2$ holes. [^12] They assume that the associated category is rigid. However this assumption is superfluous thanks to the resent result of Etingof and Penneys [14]. These relations will also be satisfied in a $\chi$-truncated modular functor for $\chi\leq-3$, as these surfaces are accessible. Hence their proof also implies that the ribbon fusion category associated to a $\chi$-truncated modular functor for $\chi\leq-3$ is modular.
Thanks to Theorem 5.9, it remains only to prove that $\mathfrak{mod}_{c}^{\geq\chi}\longrightarrow\mathfrak{Mod}$, $\mathfrak{mod}^{0,\geq\chi}\longrightarrow\mathfrak{Rib}$, $\mathfrak{ribfun}^{0,\geq\chi}\longrightarrow\mathfrak{Rib}$, and $\mathfrak{braidfun}^{\geq\chi}\longrightarrow\mathfrak{Braid}$ are fully-faithful. The point is to notice that the proofs given for $\chi=-\infty$ in the next subsections apply directly to the truncated case. Let us comment on the modular case, as the other cases are easier. In that case, the only non-trivial part of the proof is the use of curve operators, which requires $2$ points (see Section 5.6): that Dehn twists be polynomials in the curve operators and that $g=1$, $n=3$ be accessible. The first point is a direct consequence of modularity of the associated ribbon fusion category. As for the second point, $\chi\leq-3$ ensures that $g=1$, $n=3$ is accessible. ∎
5.4. Fullness for braided and ribbon functors
We prove fullness of $\mathfrak{braidfun}\rightarrow\mathfrak{Braid}$. The case of $\mathfrak{ribfun}\rightarrow\mathfrak{Rib}$ is similar, see Remark 5.13 below.
Let $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ be $2$ braided functors with respective braided fusion categories $\mathrm{C}_{\mathcal{V}_{1}}$ and $\mathrm{C}_{\mathcal{V}_{2}}$. Let $G:\mathrm{C}_{\mathcal{V}_{1}}\rightarrow\mathrm{C}_{\mathcal{V}_{2}}$ be a morphism. We may assume, without loss of generality, that the two functors have the same set of colors $\Lambda$ and that $G$ acts trivially on $\Lambda$. For $D^{n}$ the standard disk (see Section 5.1) and $\mu,\lambda_{1},\dotsc,\lambda_{n}\in\Lambda$, Proposition 5.4 provides canonical isomorphisms
$\mathcal{V}_{i}(D^{n},\mu;\lambda_{1},\dotsc,\lambda_{n})\simeq\hom_{\mathrm{C}_{\mathcal{V}_{i}}}([\mu],[\lambda_{1}]\otimes\dotsb\otimes[\lambda_{n}])$
for $i=1,2$, which are compatible with braid group actions and with the standard gluing of standard disks described in Definition 5.1 and LABEL:figurediskgluing. Hence $G$ provides isomorphisms
$G_{n,\mu,\underline{\lambda}}:\mathcal{V}_{1}(D^{n},\mu;\lambda_{1},\dotsc,\lambda_{n})\simeq\mathcal{V}_{2}(D^{n},\mu;\lambda_{1},\dotsc,\lambda_{n})$
compatible with standard gluings and braid group actions.
To show fullness, we need to extend them to any configuration of disks. Fix $D$ a $\Lambda$-labeled configuration of $n$ disks for some $n\geq 0$. Then there exists an isomorphism $\phi:D\simeq(D^{n},\mu;\lambda_{1},\dotsc,\lambda_{n})$ for some choice of $\mu,\lambda_{1},\dotsc,\lambda_{n}$. We define an isomorphism $G_{D,\phi}:\mathcal{V}_{1}(D)\simeq\mathcal{V}_{2}(D)$ by imposing the commutation of
[…]
In fact $G_{D,\phi}$ does not depend on $\phi$ or the choice of the ordering of the $\lambda_{i}$. Indeed, if $\phi^{\prime}:D\simeq(D^{n},\mu;\lambda_{1}^{\prime},\dotsc,\lambda_{n}^{\prime})$ is another isomorphism, then there exists a unique isomorphism $\alpha:(D^{n},\mu;\lambda_{1},\dotsc,\lambda_{n})\simeq(D^{n},\mu;\lambda_{1}^{\prime},\dotsc,\lambda_{n}^{\prime})$ compatible with $\phi$ and $\phi^{\prime}$. As $\alpha$ is a braid, $G_{n,\mu,\underline{\lambda}}$ and $G_{n,\mu,\underline{\lambda}^{\prime}}$ are compatible with it, and thus we have commutation of
[…]
Remark 5.13.
The proof works as is in the case of ribbon functors, replacing braid groups by framed braid groups and $\mathrm{Disk}_{\Lambda}^{b}$ by $\mathrm{Disk}_{\Lambda}$.